OpenStudy (anonymous):

10 in. = ______ feet (give the answer in a fraction)

OpenStudy (anonymous):

5/6

OpenStudy (anonymous):

how to show my work

OpenStudy (anonymous):

 you simplify the fraction

hero (hero):

12 in = 1 ft 1 in = (1/12) ft 10 in = x ft 10 (1/12) ft = x ft 10(1/12) ft/ft = x 10/12 = x 5/6 = x